# Implicit High-Order Moment Tensor Estimation and Learning Latent Variable Models

### v1

@@ Abstract
We study the task of learning latent-variable models on $\mathbb{R}^{d}$ with $k$ hidden parameters. A common obstacle towards designing computationally efficient algorithms for such models is the necessity of approximating moment tensors of super-constant degree, scaling with the parameter $k$ or the desired accuracy $\epsilon>0$. Motivated by such learning applications, we develop a general efficient algorithm for implicit moment tensor computation. Roughly speaking, our algorithm computes in $\mathrm{poly}(d,k)$ time a succinct approximate description of tensors of the form $M_{m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}$, for $w_{i}\in\mathbb{R}_{+}$—even for $m=\omega(1)$—assuming that there exists a polynomial-size arithmetic circuit whose expected output on an appropriate samplable distribution is equal to $M_{m}$, and whose covariance on this input is bounded. Our framework broadly generalizes, both conceptually and technically, the work of [LL22] which developed an efficient algorithm for the specific moment tensors that arise in the task of clustering mixtures of spherical Gaussians.
By leveraging our general implicit moment estimation algorithm, we obtain the first $\mathrm{poly}(d,k)$-time learning algorithms for the following classical latent-variable models—thereby resolving or making significant progress towards a number of important open problems in the literature.
@@
  • Learning Mixtures of Spherical Gaussians Given i.i.d. samples from a $k$-mixture of identity covariance Gaussians on $\mathbb{R}^{d}$, the goal is to learn the target mixture. We give a $\mathrm{poly}(d,k,1/\epsilon)$-time density estimation algorithm, where $\epsilon$ is the desired total variation error, under the condition that the means lie in a ball of radius $O(\sqrt{\log k})$. Prior algorithms incur super-polynomial complexity in $k$. For parameter estimation, we give a $\mathrm{poly}(d,k,1/\epsilon)$-time algorithm where $\epsilon$ is the target accuracy, under the optimal mean separation of $\Omega(\log^{1/2}(k/\epsilon))$ and the condition that the largest distance is comparable to the smallest. Prior polynomial-time parameter estimation algorithms require separation $\Omega(\log^{1/2+c}(k/\epsilon))$, for $c>0$.
  • PAC Learning Positive Linear Combinations of ReLUs Given labeled examples $(x,y)$ with $x\sim N(0,I)$ and $y=F(x)$, where $F$ is an unknown positive linear combination of $k$ ReLUs (or other reasonable activations) on $\mathbb{R}^{d}$, the goal is to PAC learn the target function in $L_{2}$-norm. We give an algorithm for this task with complexity $\mathrm{poly}(d,k)2^{\mathrm{poly}(1/\epsilon)}$, where $\epsilon$ is the desired relative $L_{2}$-error. Our is the first algorithm for this learning task that runs in fully-polynomial (i.e., $\mathrm{poly}(d,k)$) time for subconstant values of the error $\epsilon=o_{k,d}(1)$. All previously known algorithms have super-polynomial complexity, for any $\epsilon=o_{k,d}(1)$.
@@ 1 Introduction
Here we focus on the regime where both the dimension $d$ and the number of hidden parameters $k$ are large. In this context, we are interested in designing learning algorithms for natural models with complexity $\mathrm{poly}(d,k)$, i.e., polynomial in both $d$ and $k$. Of course, there are examples where such a complexity guarantee may not be possible, due to fundamental computational limitations. One such example is the task of learning (either proper learning or density estimation) $k$-mixtures of Gaussians with non-spherical covariances on $\mathbb{R}^{d}$, for which recent work [DKS17, BRST21, GVV22, DKPP24] has provided formal evidence of computational hardness—ruling out $\mathrm{poly}(d,k)$ runtime, even though $\mathrm{poly}(d,k)$ samples suffice. For the latent-variable models we study in this paper—namely $k$-mixtures of linear regressions, $k$-mixtures of spherical Gaussians, and one-hidden-layer neural networks with $k$ gates— there is no known formal evidence of hardness. On the other hand, the best known algorithms prior to this work had super-polynomial complexity (as a function of $k$). As a corollary of our main result (Proposition 1.7), we design the first $\mathrm{poly}(d,k)$ time learning algorithms for all of these problems. We thus resolve, or make significant progress towards, several well-known open problems in the learning theory literature. All of our algorithmic results rely on a novel methodology that is of broader interest and may find other applications.
Before we describe our main technique, we start by highlighting our algorithmic results on learning latent-variable models.
@@ 1.1 Efficiently Learning Latent Variable Models
Here we give the first polynomial-time algorithm. Specifically, we show (see Theorem 4.19):
@@
Let $F$ be a $k$-MLR distribution on $\mathbb{R}^{d+1}$ with $B,\sigma\leq 1$, where $B=\max_{i}\|\beta_{i}\|_{2}$ and $\sigma$ is the standard deviation of the Gaussian noise. There exists an algorithm that draws $N=\mathrm{poly}(k,d)(1/\epsilon)^{O(\sigma^{-2})}$ samples from $F$, runs in $\mathrm{poly}(N,d)$ time, and returns a sampler for a distribution that is $\epsilon$-close to $F$ in total variation distance.
The existence of a learning algorithm for $k$-mixtures of linear regressions with a $\mathrm{poly}(d,k)$ complexity was a well-known open problem in the literature. This question was most recently posed as one of the main open problems in [CLS20]; see the associated talk at STOC 2020 [Che20]. Theorem 1.1 answers this open question in the affirmative.
It should also be noted that the requirement that $\sigma,B\leq 1$ can be removed at the cost of making the runtime $\mathrm{poly}(k,d)(1/\epsilon)^{1+O((B/\sigma)^{2})}$. This is done by computing an upper bound $A$ on $B+\sigma$ (for example by taking many samples and returning something proportional to the largest value of $|y|$), and renormalizing by replacing $y$ by $y/A$. An interesting open question that remains is to remove the dependence on $\sigma,B$ from the complexity of our algorithm.
@@
Learning One-hidden-layer ReLU Networks
A one-hidden-layer ReLU network is a function $F:\mathbb{R}^{d}\to\mathbb{R}$ of the form $F(x)=\sum_{i=1}^{k}w_{i}\mathrm{ReLU}(v_{i}\cdot x)$ for some $w_{i}\in\mathbb{R}$ and unit vectors $v_{i}\in\mathbb{R}^{d}$, where $\mathrm{ReLU}:\mathbb{R}\to\mathbb{R}$ is defined as $\mathrm{ReLU}(t)\stackrel{{\scriptstyle{\mathrm{\footnotesize def}}}}{{=}}\max\{0,t\}$ [^2] Our result does not require the assumption that the activation function is a ReLU. The only properties required is that the activation has bounded higher moments and non-vanishing even-degree Fourier coefficients. Our algorithm can be straightforwardly extended to activation functions satisfying these properties.. Here we focus on the natural setting that the weights $w_{i}$ are positive. Let $\mathcal{C}_{d,k}$ be the corresponding class of functions. The PAC learning problem for $\mathcal{C}_{d,k}$ is the following: The input is a multiset of i.i.d. labeled examples $(x,y)$, where $x\sim N(0,I)$ and $y=F(x)$, for some $F\in\mathcal{C}_{d,k}$. The goal is to output a hypothesis $H:\mathbb{R}^{d}\to\mathbb{R}$ that with high probability is close to $F$ in $L_{2}$-norm.
A central question in the foundations of deep learning is whether this class (sometimes termed “sums of ReLUs”) is learnable in polynomial time, when the marginal distribution is Gaussian. Quoting [CKM21]:
@@
Prior to our work, the fastest known algorithm [DK20] had super-polynomial complexity, even for constant accuracy. Our main result for this problem is the following (see Theorem 4.7):
Theorem 1.4 (PAC Learning $\mathcal{C}_{d,k}$).
There is an algorithm that given $\epsilon>0$, and access to examples $(x,F(x))$, where $x\sim N(0,I)$ and $F\in\mathcal{C}_{d,k}$ is the target function, it runs in time $\mathrm{poly}(d,k)2^{\mathrm{poly}(1/\epsilon)}$, and outputs a hypothesis $H:\mathbb{R}^{d}\to\mathbb{R}$ such that with high probability $\|H-F\|_{2}\leq\epsilon\|F\|_{2}$.
Theorem 1.4 makes significant progress towards resolving the aforementioned open problem. Specifically, it gives the first learning algorithm for sums of ReLUs that runs in $\mathrm{poly}(d,k)$ time for sub-constant values of $\epsilon=o(1)$—namely up to $\epsilon=1/\log^{c}(dk)$, for some constant $c>0$. The fastest previous algorithm [DK20] has sample and computational complexity $\mathrm{poly}(d)(k/\epsilon)^{O(\log^{2}(k))}$, i.e., quasi-polynomial in $k$ even for $\epsilon=\Theta(1)$. On the other hand, the running time of the [DK20] algorithm is better than ours for $\epsilon$ very small (e.g., $1/\mathrm{poly}(d,k)$).
@@
For one-hidden-layer ReLU networks with general (i.e., positive or negative) weights, the fastest known algorithms take time $(d/\epsilon)^{\mathrm{poly}(k)}$ [CN24, DK24]. Moreover, cryptographic hardness results [LZZ24] rule out the existence of an algorithm with $\mathrm{poly}(d,k)$ runtime.
@@ 1.2 Main Result: General Implicit Moment Tensor Estimation
All aforementioned learning results are obtained as applications of a general implicit moment tensor computation algorithm that we design, which is the focus of this section. We start by describing our framework, under which such an algorithm is possible, followed by the statement of our main result (Proposition 1.7) and a comparison with the prior work [LL22].
Sequential Tensor Computations and Implicit Moment Tensor Estimation
@@
Roughly speaking, our main result shows that we can efficiently compute a succinct approximate description of tensors of the form $M_{m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}$, for non-negative $w_{i}$’s, even for large (aka super-constant) values of the order parameter $m$, assuming there exists a polynomial-size arithmetic circuit whose expected output—under an appropriate samplable distribution—is equal to $M_{m}$ and whose covariance (under the same distribution) is bounded above.
To formally state our result, we introduce the following key notion:
Definition 1.6 (Sequential Tensor Computation).
  • (i)
    Multiplication of a tensor, vector, or scalar by a scalar.
  • (ii)
  • (iii)
We note that an STC is a general way to produce higher-order tensors that is compatible with the kinds of computations that we need to perform on them (see Lemmas 3.4 and 3.5) in the context of learning applications. Our main algorithmic result is the following:
Proposition 1.7 (Main Algorithmic Result).
@@
Let $w_{1},\ldots,w_{k}\in\mathbb{R}_{\geq 0}$ and $v_{1},\ldots,v_{k}\in\mathbb{R}^{d}$. Consider the sequence of dimension-$d$ tensors $\{M_{t}\}$ with $M_{t}$ order-$t$, for $t\in\mathbb{Z}_{+}$, defined by $M_{t}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}$.
Let $m\in\mathbb{Z}_{+}$. Suppose that for all positive integers $t\leq 2m$, with $t$ even or equal to $m$, there exists an STC $\mathcal{S}_{t}$ of order-$t$ and size at most $S$ with the following property: when given as input $\mathcal{I}$ a sample drawn from some efficiently samplable distribution $\mathcal{D}_{t}$, the output tensor $\mathcal{S}_{t}(\mathcal{I})$, $\mathcal{I}\sim\mathcal{D}_{t}$, has mean $M_{t}$ and covariance bounded above by $V$, for some $V>0$. Let $\mathcal{F}_{m}$ be another STC of order-$m$ and size at most $S$, whose input is partitioned as $(\mathcal{X},\mathcal{Y})$, with the following properties: when $\mathcal{X}\sim X$ and $\mathcal{Y}\sim\mathcal{D}^{\prime}$, where $\mathcal{D^{\prime}}$ is efficiently samplable, the tensor $\mathcal{F}_{m}(\mathcal{X},\mathcal{Y})$ has $\mathbf{E}_{\mathcal{Y}\sim\mathcal{D^{\prime}}}[\mathcal{F}_{m}(\mathcal{X},\mathcal{Y})]=T(\mathcal{X})$ and second moment over both $\mathcal{X}$ and $\mathcal{Y}$ bounded by $V$.
$\mathbf{E}_{\mathcal{X}\sim X}[(\mathcal{A}-\langle T(\mathcal{X}),M_{m}\rangle)^{2}]\leq\frac{\mathrm{poly}\left(k,m,d,V,1/\tau,1+\mathop{\textstyle\sum}_{i=1}^{k}w_{i}\right)\left(1+\max_{i\in[k]}\|v_{i}\|_{2}\right)^{2m}}{N^{1/2}}\;.$
In words, we show that, with high probability over the samples drawn from $\mathcal{D}_{t}$ and $\mathcal{D}^{\prime}$, the expected squared error of our estimator over a random choice of $\mathcal{X}\sim X$ is bounded as shown above. In particular, our algorithm can be thought of as taking samples from $\mathcal{D}_{t}$ and $\mathcal{D}^{\prime}$ as input and returning an evaluator for a function of $\mathcal{X}$ that with high probability is close in $L_{2}$ to the function $\langle T(\mathcal{X}),M_{m}\rangle$.
Some remarks are in order. First, we did not attempt to optimize the polynomial dependence on the relevant parameters, as this would not affect the qualitative aspects of our learning theory applications. Second, the dependence on the failure probability $\tau$ can be made logarithmic, i.e., $\log(1/\tau)$, by running the algorithm $O(\log(1/\tau))$ times and taking a median; such a dependence is not required in our applications.

### v2

@@ Abstract
We study the general task of learning latent-variable models on $\mathbb{R}^{d}$ with $k$ hidden parameters. A common technique to address this task algorithmically is (some version of) the method of moments. Unfortunately, moment-based approaches are often hampered by the fact that the moment tensors of super-constant degree cannot even be written down in polynomial time. Motivated by such learning applications, we develop a general efficient algorithm for implicit moment tensor computation. Roughly speaking, our algorithm computes in $\mathrm{poly}(d,k)$ time a succinct approximate description of tensors of the form $M_{m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}$, for $w_{i}\in\mathbb{R}_{+}$—even for $m=\omega(1)$—assuming that there exists an unbiased estimator for $M_{m}$ with small variance that takes an appropriately nice form. Our framework broadly generalizes, both conceptually and technically, the work of [LL22] which developed an efficient algorithm for the specific moment tensors that arise in the task of clustering mixtures of spherical Gaussians.
By leveraging our implicit moment estimation algorithm, we obtain the first $\mathrm{poly}(d,k)$-time learning algorithms for the following classical latent-variable models—thereby resolving or making significant progress towards a number of important open problems in the literature.
@@
  • Mixtures of Spherical Gaussians Given i.i.d. samples from a $k$-mixture of identity covariance Gaussians on $\mathbb{R}^{d}$, the goal is to learn the target mixture. For density estimation, we give a $\mathrm{poly}(d,k,1/\epsilon)$-time learning algorithm, where $\epsilon$ is the desired total variation error, under the condition that the means lie in a ball of radius $O(\sqrt{\log k})$. Prior algorithms incur super-polynomial complexity in $k$. For parameter estimation, we give a $\mathrm{poly}(d,k,1/\epsilon)$-time algorithm where $\epsilon$ is the target accuracy, under the optimal mean separation of $\Omega(\log^{1/2}(k/\epsilon))$ and the condition that the largest distance is comparable to the smallest. Prior polynomial-time parameter estimation algorithms require separation $\Omega(\log^{1/2+c}(k/\epsilon))$, for $c>0$.
  • Positive Linear Combinations of Non-Linear Activations Given i.i.d. samples $(x,y)$ with $x\sim N(0,I)$ and $y=F(x)$, where $F$ is a positive linear combination of $k$ reasonable non-linear activations on $\mathbb{R}^{d}$, the goal is to learn the target function in $L_{2}$-norm. Our main result is a general algorithm for this task with complexity $\mathrm{poly}(d,k)g(\epsilon)$, where $\epsilon$ is the desired error and the function $g$ depends on the Hermite concentration of the target class of functions. Specifically, for positive linear combinations of ReLU activations, our algorithm has complexity $\mathrm{poly}(d,k)2^{\mathrm{poly}(1/\epsilon)}$. This is the first algorithm for this class that runs in $\mathrm{poly}(d,k)$ time for sub-constant values of $\epsilon=o_{k,d}(1)$. Finally, for positive linear combinations of cosine activations with bounded frequency, our algorithm runs in $\mathrm{poly}(d,k,1/\epsilon)$ time.
@@ 1 Introduction
Here we focus on the regime where both the dimension $d$ and the number of hidden parameters $k$ are large. In this context, we are interested in designing learning algorithms for natural models with complexity $\mathrm{poly}(d,k)$, i.e., polynomial in both $d$ and $k$. Of course, there are examples where such a complexity upper bound may not be possible, due to fundamental computational limitations. A prominent such example (where $\mathrm{poly}(d,k)$ complexity is unlikely) is for the task of learning (proper learning or density estimation) $k$-mixtures of Gaussians on $\mathbb{R}^{d}$ with general—aka non-spherical—covariances. For this learning task, recent work [DKS17, BRST21, GVV22, DKPP24] has provided strong evidence of computational hardness—ruling out $\mathrm{poly}(d,k)$ runtime, even though $\mathrm{poly}(d,k)$ samples information-theoretically suffice.
For the latent-variable models we study in this paper—namely $k$-mixtures of linear regressions, $k$-mixtures of spherical Gaussians, and one-hidden-layer neural networks with $k$ gates— there is no known formal evidence of hardness. On the other hand, prior to this work, the best known algorithms had super-polynomial complexity (as a function of the parameter $k$). By leveraging our main result (Proposition 1.8), we design the first $\mathrm{poly}(d,k)$ time learning algorithms for all of these problems. We thus resolve, or make significant progress towards, several well-known open problems in the learning theory literature. All of our learning algorithms rely on a novel methodology—which we term implicit moment tensor estimation—that is of broader interest and we believe may find other applications.
Before we describe our main technique, we start by highlighting its algorithmic implications on learning latent-variable models.
@@ 1.1 Efficiently Learning Latent Variable Models
As our main contribution for this problem, we give the first $\mathrm{poly}(d,k)$ time algorithm. Specifically, we show (see Theorem 4.22):
@@
Let $F$ be a $k$-MLR distribution on $\mathbb{R}^{d+1}$ with $B,\sigma\leq 1$, where $B=\max_{i}\|\beta_{i}\|_{2}$ and $\sigma>0$ is the standard deviation of the Gaussian noise. There exists an algorithm that draws $N=\mathrm{poly}(k,d)(1/\epsilon)^{O(\sigma^{-2})}$ samples from $F$, runs in $\mathrm{poly}(N,d)$ time, and returns a sampler for a distribution that is $\epsilon$-close to $F$ in total variation distance.
We note that the existence of a learning algorithm for $k$-MLR with a $\mathrm{poly}(d,k)$ complexity had been a well-known open problem in the literature. This question was most recently posed as one of the main open problems in [CLS20]; see the associated talk at STOC 2020 [Che20]. Theorem 1.1 answers this open question in the affirmative.
Some additional remarks are in order. For the setting that the standard deviation $\sigma$ is at least a positive universal constant, the complexity of our density estimation algorithm is $\mathrm{poly}(k,d,1/\epsilon)$—essentially resolving the complexity of this learning task. If $\sigma$ is very small (but still non-zero), for example for $\sigma=\Theta(\epsilon)$ (a regime considered in [CLS20]), the complexity of our learning algorithm becomes $\mathrm{poly}(k,d)2^{\tilde{O}(1/\epsilon^{2})}$. Even in this regime, we obtain the first learning algorithm for $k$-MLR that runs in $\mathrm{poly}(d,k)$ time for sub-constant values of $\epsilon=o(1)$—namely up to $\epsilon=1/\log^{1/2}(dk)$. The fastest previous algorithm [DK20] has complexity quasi-polynomial in $k$ even for constant $\epsilon$. An interesting question left open by our work is to remove the dependence on $\sigma$ from the complexity of our algorithm.
@@
Learning One-hidden-layer Neural Networks
A one-hidden-layer neural network is any function $F:\mathbb{R}^{d}\to\mathbb{R}$ of the form $F(x)=\sum_{i=1}^{k}w_{i}\sigma(v_{i}\cdot x)$ for some $w_{i}\in\mathbb{R}$ and unit vectors $v_{i}\in\mathbb{R}^{d}$, where $\sigma:\mathbb{R}\to\mathbb{R}$ is a known activation. In this paper, we focus on the setting that the weights $w_{i}$ are positive (for general weights, recent work has given evidence of computational hardness ruling out $\mathrm{poly}(d,k)$ runtime; see Remark 1.7).
Let $\mathcal{C}_{\sigma,d,k}$ be the corresponding class of functions. The PAC learning problem for $\mathcal{C}_{\sigma,d,k}$ is the following: The input is a multiset of i.i.d. labeled examples $(x,y)$, where $x\sim N(0,I)$ and $y=F(x)$, for some $F\in\mathcal{C}_{\sigma,d,k}$, where the activation $\sigma$ is known to the algorithm. The goal is to output a hypothesis $H:\mathbb{R}^{d}\to\mathbb{R}$ that with high probability is close to $F$ in $L_{2}$-norm.
A prominent special case corresponds to $\sigma(u)=\mathrm{ReLU}(u)\stackrel{{\scriptstyle{\mathrm{\footnotesize def}}}}{{=}}\max\{0,u\}$. The learnability of one-hidden-layer ReLU networks has been extensively studied over the past decade; see, e.g., [GLM18, BJW19, DKKZ20, DK20, CKM21, CDG+23, CN24a, DK24].
A central question in the foundations of deep learning is whether this class is learnable in polynomial time, when the marginal distribution is Gaussian. Quoting [CKM21] (see also Chen’s thesis [Che21]):
@@
We give a general algorithm for $\mathcal{C}_{\sigma,d,k}$ that succeeds for a wide class of activations $\sigma$, including ReLUs and cosine activations. Roughly speaking, the only properties required on $\sigma$ is that it has bounded fourth moment and non-vanishing even-degree Hermite coefficients. Specifically, we prove the following (see Theorem 4.7):
Theorem 1.4 (PAC Learning $\mathcal{C}_{\sigma,d,k}$).
Suppose that the activation $\sigma$ has $\|\sigma\|_{4}=O(1)$. Letting $c_{\sigma,t}:=\mathbf{E}[\sigma(G)h_{t}(G)]$ be the $t^{th}$ Hermite coefficient of $\sigma$, suppose additionally that for some $\epsilon,\delta>0$ and some $n\in\mathbb{Z}_{+}$ it holds: (i) $\sum_{t>n}c_{\sigma,t}^{2}<\epsilon^{2}/4$, and (ii) $|c_{\sigma,t}|\geq\delta$ for all $1\leq t\leq 2n$ unless $t$ is odd and $c_{\sigma,t}=0$.
Then there is an algorithm that given $\epsilon$, $\sigma,k,d$, and access to examples $(x,F(x))$, where $x\sim N(0,I)$ and $F\in\mathcal{C}_{\sigma,d,k}$, it runs in time $\mathrm{poly}(dk/(\epsilon\delta))2^{O(n)}$, and outputs a hypothesis $H:\mathbb{R}^{d}\to\mathbb{R}$ such that with high probability $\|H-F\|_{2}\leq\epsilon$.
For the special case where $\sigma(u)=\mathrm{ReLU}(u)$, this leads to the following (see also Corollary 4.9):
Corollary 1.5.
There is a PAC learning algorithm for positive linear combinations of $k$ ReLUs on $\mathbb{R}^{d}$ with complexity $\mathrm{poly}(d,k)2^{\mathrm{poly}(1/\epsilon)}$.
Corollary 1.5 makes significant progress towards resolving the above open problem [CKM21, Che21]. Specifically, it gives the first learning algorithm for sums of ReLUs that runs in $\mathrm{poly}(d,k)$ time for sub-constant values of $\epsilon=o(1)$—namely up to $\epsilon=1/\log^{c}(dk)$, for some constant $c>0$. The fastest previous algorithm [DK20] has sample and computational complexity $\mathrm{poly}(d)(k/\epsilon)^{O(\log^{2}(k))}$, i.e., quasi-polynomial in $k$ even for $\epsilon=\Theta(1)$. On the other hand, the running time of the [DK20] algorithm is better than ours for $\epsilon$ very small (e.g., $1/\mathrm{poly}(d,k)$).
Our second application is for positive linear combinations of periodic activations, in particular cosine activations. Such periodic activations are of particular interest in signal processing and computer vision, where they have been empirically observed to be more accurate that ReLU networks in representing natural signals; see, e.g., [SMB+20, MST+20, RNM+21, VPHA24].
For $\sigma(u)=\cos(\gamma u)$, where $\gamma>0$ is a parameter, we obtain the following (see Corollary 4.10):
Corollary 1.6.
For $\sigma(u)=\cos(\gamma u)$, for some frequency parameter $\gamma>0$, there is a PAC learning algorithm for $\mathcal{C}_{\sigma,d,k}$ with complexity $2^{O(1/\gamma^{2})}\mathrm{poly}(dk/\epsilon)$.
While the dependence on $\gamma$ in our algorithm is exponential, this is inherent due to matching computational lower bounds [SVWX17, SZB21, DKRS23] (even for $k=1$). Note that for $\gamma\ll 1/\sqrt{\log(dk/\epsilon)}$, our algorithm runs in $\mathrm{poly}(dk/\epsilon)$ time.
@@
For one-hidden-layer ReLU networks with general (i.e., positive or negative) weights, the fastest known algorithms take time $(d/\epsilon)^{\mathrm{poly}(k)}$ [CN24b, DK24]. Moreover, cryptographic hardness results [LZZ24] rule out the existence of an algorithm with $\mathrm{poly}(d,k)$ runtime.
@@ 1.2 Main Result: General Implicit Moment Tensor Estimation
All aforementioned learning results are obtained as applications of a general implicit moment tensor computation algorithm that we design, which is the focus of this section. We start by describing our framework, under which such an algorithm is possible, followed by an informal statement of our main result (Proposition 1.8) and a comparison with the most relevant prior work [LL22].
Background on Tensors
In order to proceed, we require basic background on tensors (see Section 2 for more details). For a vector $x\in\mathbb{R}^{d}$, we denote by $x^{\otimes m}$ the $m$-th order tensor power of $x$. Let $\otimes$ denote the tensor/Kronecker product between vector spaces. Fix some $t\in\mathbb{Z}_{+}$. For vector spaces $V_{i}$, $i\in[t]$, an order-$t$ tensor is an element $A\in\bigotimes_{i=1}^{t}V_{i}$. If $A,B\in\bigotimes_{i=1}^{t}V_{i}$, for some inner product spaces $V_{i}$, $i\in[t]$, then we use $\langle A,B\rangle$ to denote the inner product of $A$ and $B$ induced by the inner products on $V_{i}$. Namely, using the unique inner product structure on $\bigotimes_{i=1}^{t}V_{i}$ so that $\langle v_{1}\otimes\cdots\otimes v_{t},w_{1}\otimes\cdots\otimes w_{t}\rangle=\prod_{i=1}^{t}\langle v_{i},w_{i}\rangle$ for all vectors $v_{i},w_{i}\in V_{i}$. For the special case that $V_{i}=\mathbb{R}^{d_{i}}$, $d_{i}\in\mathbb{Z}_{+}$, a tensor $A$ is defined by a $t$-dimensional array with real entries $A_{\alpha}$, where $\alpha=(\alpha_{1},\ldots,\alpha_{t})$ with $\alpha_{i}\in[d_{i}]$. For tensors $A,B$ of order $t$ with the same dimensions, the inner product $\langle A,B\rangle$ is the entrywise inner product of the corresponding arrays.
Implicit Moment Tensor Estimation
@@
Roughly speaking, our algorithmic result shows that it is possible to efficiently compute a succinct approximate description of tensors of the form $M_{m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}$, for non-negative $w_{i}$’s, even for large (aka super-constant) values of the order parameter $m$, assuming there exists a nice polynomial-size arithmetic circuit whose expected output—under an appropriate sampleable distribution—is equal to $M_{m}$ and whose covariance (under the same distribution) is not too large.
Having a succinct description of a higher order tensor begs the question of what exactly we can do with this description; as we cannot hope, for example, to efficiently decompress the description back into the full tensor. Instead, we could hope that for the following: for some reasonable class of functions $f:(\mathbb{R}^{d})^{\otimes m}\to\mathbb{R}$, our succinct description allows us to efficiently compute approximations to $f(M_{m})$. A natural class of such functions involves taking the inner product of $M_{m}$ with some tensor $T$ (of the same order and dimension). Of course, to have a hope of achieving this, $T$ itself will need to have a concise representation—which for our purposes means that $T$ is the expected output of a nice arithmetic circuit on some sampleable random input distribution. Fortunately, this limited means of querying $M_{m}$ turns out to be sufficient for many learning theoretic applications. For further details on these applications, see Section 1.3.2.
We are now ready to state an informal description of our algorithmic result in this context (see Definition 3.1 and Proposition 3.2 for the detailed formal versions).
Proposition 1.8 (Implicit Moment Tensor Computation, Informal).
@@
Let $d,k,m\in\mathbb{Z}_{+}$, $w_{i}\in\mathbb{R}_{\geq 0}$ and $v_{i}\in\mathbb{R}^{d}$ for all $i\in[k]$. Consider the sequence of dimension-$d$ tensors $\{M_{t}\}$ with $M_{t}$ order-$t$, for $t\in\mathbb{Z}_{+}$, defined by $M_{t}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}$. Suppose that both assumptions below hold:
  1. 1.
    For all positive integers $t\leq 2m$, with $t$ even or equal to $m$, we have a reasonable way to sample from a distribution $\mathcal{S}_{t}$ over order-$t$ tensors that has expectation $M_{t}$ and variance at most $V$.
  2. 2.
    We have a reasonable way to sample from a distribution $\mathcal{F}_{m}$ over order-$m$ tensors that has expectation $T$ and variance at most $V$.
Then there is an algorithm that draws $N$ such samples, runs in time $\mathrm{poly}(N,d,k,m)$, and computes an approximation to $\langle T,M_{m}\rangle$ that with high probability has absolute error
$\frac{\mathrm{poly}\left(k,m,d,V\right)\left(1+\max_{i\in[k]}\|v_{i}\|_{2}\right)^{m}}{N^{1/4}}\;.$
Discussion
Some comments are in order. Recall that the goal is to efficiently compute a succinct approximation to $M_{m}$ for a given $m\in\mathbb{Z}_{+}$. Formally speaking, the distribution over order-$t$ tensors (with expected value $M_{t}$) that we are assumed to have sample access to is the output of a nice arithmetic circuit (in particular, what we term a Sequential Tensor Computation; see Definition 3.1) on some sampleable distribution. The algorithm of Proposition 1.8 outputs an approximation of the inner product $\langle T,M_{m}\rangle$, for a tensor $T$ that is the expected output of a nice arithmetic circuit on some sampleable distribution.